# Non-minimally Coupled Running Curvaton: A Unified Approach to Early-Universe Inflation and Phantom Dark Energy

### v1

@@ I Introduction
This observational hint poses a significant theoretical challenge. Canonical single-field quintessence models are typically restricted to the region $w\geq-1$ Caldwell et al. (1998); Zlatev et al. (1999), while naive attempts to realize phantom behavior often lead to pathologies such as ghost degrees of freedom or gradient instabilities Vikman (2005); Hu (2005); Caldwell et al. (2003). Consequently, constructing a consistent theoretical framework that explains both the early inflationary epoch and the specific late-time phantom dynamics favored by DESI—without introducing instabilities—remains an open problem Creminelli et al. (2009); Feng et al. (2005); Nojiri et al. (2005). However, this issue can be addressed by introducing a non-minimal coupling Yao et al. (2025); Wolf et al. (2026); Tiwari et al. (2025); Sánchez López et al. (2025); Wang et al. (2025); Tita et al. (2025); Wolf et al. (2025); Adam et al. (2025). Although numerous approaches exist to study evolving dark energy—including the Quintom scenario Cai et al. (2025); Yang et al. (2025, 2024), we restrict our attention in this paper to the non-minimally coupled dark energy model.
@@ IV.3 Modified Friedmann Equations
In a flat FLRW universe, the variation of Action (1) yields the modified Friedmann equations. Neglecting the inflaton field at late times, we have:
@@ IV.6 Constraints from DESI
To constrain the free parameters of the non-minimal coupling model, namely $\theta=\{\xi,\lambda,V_{0},V_{1}\}$, we performed a Monte Carlo parameter scan using the rejection sampling technique. We assumed uniform priors for all parameters within physically motivated ranges: $\xi\in[0,5]$, $\lambda\in[0,2]$, $V_{0}\in[0.5,1.5]$, and $V_{1}\in[1.0,6.0]$. For each randomly sampled set of parameters, we numerically solved the background equations of motion (Eq. 32 and Eq. 33) to derive the corresponding dark energy equation-of-state parameters at present day, $w_{0}$, and its evolution, $w_{a}$. These theoretical predictions were then mapped onto the observational $w_{0}-w_{a}$ plane. We utilized the $95\%$ confidence level (C.L.) contours obtained from the combination of DESI DR2 BAO, Pantheon+ supernovae, and Planck CMB data as our acceptance criterion. Specifically, we adopted a top-hat likelihood approach: a parameter set was accepted only if its predicted $(w_{0},w_{a})$ fell within the observational $95\%$ C.L. region; otherwise, it was rejected. Fig. 2 shows the resulting marginalized posterior distributions and 2D confidence regions for the model parameters, derived from $20,000$ Monte Carlo trials. From this figure, the best fit parameter $(\xi,\lambda,V_{0},V_{1})=(2.9,0.96,0.76,3.5)$. We can see the model $B$ will almost lie out the DESI constraint since the $\xi$ is smaller than the minimal value of $\xi=1.6$ in 2. Generally, Fig. 3 demonstrates the consistency with fig. 2.
@@ V Discussion and Conclusion
In summary, the Non-minimally Coupled Running Curvaton offers a theoretically sound and observationally consistent pathway to unify the physics of the early and late universe. It not only alleviates the current tensions in the dark energy sector but also provides a concrete mechanism for dynamical phantom crossing without violating fundamental stability conditions. Future high-precision surveys, such as Euclid and LSST, will further constrain the derivative of the equation of state, offering a decisive test for the specific $w_{0}-w_{a}$ trajectory predicted by this framework. Furthermore, we could extend our work into the multi-field framework Liu and Prokopec (2021); Zhang et al. (2023) and the framework of $f(R)$ gravity Liu (2018); Liu et al. (2018, 2021). The running curvaton model could also generate the primordial Black-hole (PBH) in order to account for the dark matter Liu (2023), even we could discuss the non-minimally coupling effect to the formation of PBH.
@@ C.4 Consistency Check: Sub-dominance Condition
Furthermore, the constraints from local gravity tests (Eq. 31 in the main text) imply $|\xi\chi_{0}/M_{P}|\lesssim 10^{-3}$. Assuming the field does not traverse trans-Planckian distances from inflation to today, the correction term $\xi(\chi/M_{P})^{2}$ remains negligible throughout cosmic history.

### v2

@@ I Introduction
This observational hint poses a significant theoretical challenge. Canonical single-field quintessence models are typically restricted to the region $w\geq-1$ Caldwell et al. (1998); Zlatev et al. (1999), while naive attempts to realize phantom behavior often lead to pathologies such as ghost degrees of freedom or gradient instabilities Vikman (2005); Hu (2005); Caldwell et al. (2003). Consequently, constructing a consistent theoretical framework that explains both the early inflationary epoch and the specific late-time phantom dynamics favored by DESI—without introducing instabilities—remains an open problem Creminelli et al. (2009); Feng et al. (2005); Nojiri et al. (2005). However, this issue can be addressed by introducing a non-minimal coupling Yao et al. (2025); Wolf et al. (2026); Tiwari et al. (2025); Sánchez López et al. (2025); Wang et al. (2025); Tita et al. (2025); Wolf et al. (2025); Adam et al. (2025). Although numerous approaches exist to study evolving dark energy—including the Quintom scenario Cai et al. (2025); Yang et al. (2025, 2024), $f(R)$ gravitational models Nojiri et al. (2025b, a); Odintsov and Oikonomou (2026), even using the minimally coupling to investigate the dynamical DE Gialamas et al. (2025b) that represents the local effects of DE Gialamas et al. (2025a). We restrict our attention in this paper to the non-minimally coupled dark energy model.
@@ IV.3 Modified Friedmann Equations
In a flat FLRW universe, the variation of Action (1) yields the modified Friedmann equations (Detailed calcultion can be found in Appendix D). Neglecting the inflaton field at late times, we have:
@@ IV.6 Constraints from DESI
To constrain the free parameters of the non-minimal coupling model, namely $\theta=\{\xi,\lambda,V_{0},V_{1}\}$, we performed a Monte Carlo parameter scan using the rejection sampling technique. We assumed uniform priors for all parameters within physically motivated ranges: $\xi\in[0,5]$, $\lambda\in[0,2]$, $V_{0}\in[0.5,1.5]$, and $V_{1}\in[1.0,6.0]$. For each randomly sampled set of parameters, we numerically solved the background equations of motion (Eq. 32 and Eq. 33) to derive the corresponding dark energy equation-of-state parameters at present day, $w_{0}$, and its evolution, $w_{a}$. These theoretical predictions were then mapped onto the observational $w_{0}-w_{a}$ plane. We utilized the $95\%$ confidence level (C.L.) contours obtained from the combination of DESI DR2 BAO, Pantheon+ supernovae, and Planck CMB data as our acceptance criterion. Specifically, we adopted a top-hat likelihood approach: a parameter set was accepted only if its predicted $(w_{0},w_{a})$ fell within the observational $95\%$ C.L. region; otherwise, it was rejected. Fig. 3 shows the resulting marginalized posterior distributions and 2D confidence regions for the model parameters, derived from $20,000$ Monte Carlo trials. Based on the results shown in this figure, the best-fit parameters are $(\xi,\lambda,V_{0},V_{1})=(2.9,0.96,0.76,3.5)$. We observe that Model B lies almost entirely outside the DESI constraints, as its $\xi$ value is smaller than the minimum threshold of $\xi=1.6$ shown in Fig. 2. In contrast, Model C aligns well with the best-fit region in Fig. 3. Overall, the results in Fig. 3 demonstrate strong consistency with those presented in Fig. 2.
@@ V Discussion and Conclusion
In summary, the Non-minimally Coupled Running Curvaton offers a theoretically sound and observationally consistent pathway to unify the physics of the early and late universe. It not only alleviates the current tensions in the dark energy sector but also provides a concrete mechanism for dynamical phantom crossing without violating fundamental stability conditions. Future high-precision surveys, such as Euclid and LSST, will further constrain the derivative of the equation of state, offering a decisive test for the specific $w_{0}-w_{a}$ trajectory predicted by this framework. Furthermore, this work could be extended into multi-field frameworks Liu and Prokopec (2021); Zhang et al. (2023) and $f(R)$ gravity theories Liu (2018); Liu et al. (2018, 2021). The running curvaton model also provides a mechanism for generating primordial black holes (PBHs) as candidates for dark matter Liu (2023). Additionally, one could investigate the effects of non-minimal coupling on PBH formation.
VI ACKNOWLEDGMENTS
LH, BC are supported by the National Natural Science Foundation of China (Grant No. 12165009) and the Hunan Natural Science Foundation (Grants No. 2023JJ30487 and No. 2022JJ40340).
@@ C.4 Consistency Check: Sub-dominance Condition
Assuming the field does not traverse trans-Planckian distances from inflation to today, the correction term $\xi(\chi/M_{P})^{2}$ remains negligible throughout cosmic history.